OpenStudy (anonymous):

help i will medal! Compare the pair of numbers. Use <, =, or >. 3/4 [_] 33/40

OpenStudy (anonymous):

PLZ HELP! @sammixboo @iGreen

OpenStudy (anonymous):

can you help @iGreen ?

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }\times \frac{ 10 }{ 10 }=\frac{ 30 }{ 40 }<\frac{ 33 }{ 40 }\]

OpenStudy (anonymous):

THANK YOU FOR EXPLAINING TO I AM GOING TO FAN YOU and here is your medal!!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

can you help me with one more?

OpenStudy (anonymous):

i will try.

OpenStudy (anonymous):

\[order the numbers from lest \to greatest. \frac{ 3 }{ 4 }\frac{ 19 } { 24} \frac{ 2 }{ 3 }\]

OpenStudy (anonymous):

@surjithayer you there?

OpenStudy (anonymous):

@iGreen can you help me

OpenStudy (igreen):

3/4 = 0.75 19/24 = 0.79 2/3 = 0.67 Now can you order them? @owllover123

OpenStudy (igreen):

It's gonna be: 0.67, 0.75, 0.79 or \(\dfrac{2}{3}, \dfrac{3}{4}, \dfrac{19}{24}\) @owllover123

OpenStudy (anonymous):

thank you @igreen i got it now!

OpenStudy (anonymous):

oh and by the way i fanned you!

OpenStudy (anonymous):

make the denominator same. \[\frac{ 2 }{ 3 }\times \frac{ 8 }{ 8 }=\frac{ 16 }{ 24 }\]\[\frac{ 3 }{ 4 }\times \frac{ 6 }{ 6 }=\frac{ 18 }{ 24 }\]

OpenStudy (anonymous):

\[\frac{ 16 }{ 24 }<\frac{ 18 }{ 24 }< \frac{ 19 }{ 24 }\]